Market for good X is characterized by these demand and supply schedule: p=a-bQd,
P=c-dQs ; a,b, c, and d are all positive constants
1
Expert's answer
2020-11-11T08:33:42-0500

"Qd=\frac{a-P}{b}"


"Qs=\frac{c-P}{d}"


let's say

a=300, c=200, b=200, d=100

Qd=Qs

than


"\frac{300-P}{200}=\frac{200-P}{100}"


P=100


"Qd=\frac{300-P}{200}=\frac{300-100}{200}=1"


"Qs=\frac{200-P}{100}=\frac{200-100}{200}=1"